OpenStudy (mtalhahassan2):

Need help!! 4. Solve for x: cos 3x=1/2 0

OpenStudy (mertsj):

If you want all x solutions between 0 and pi, find all 3x solutions between 0 and 3pi

OpenStudy (mertsj):

\[\cos (3x)=\frac{1}{2}\] \[3x=\frac{\pi}{3}, \frac{7\pi}{3},\frac{13\pi}{3}\]

OpenStudy (mertsj):

\[x=\frac{\pi}{9},\frac{7\pi}{9}, \frac{13\pi}{9}\]

OpenStudy (mertsj):

Whoops...one too many. Discard 13pi/9

OpenStudy (mtalhahassan2):

wait how you get 7pie/3

OpenStudy (mtalhahassan2):

I get pie/3

OpenStudy (mtalhahassan2):

know how can I find the other solutions

OpenStudy (mtalhahassan2):

Plz help me

OpenStudy (mertsj):

Add 2pi to pi/3

OpenStudy (mertsj):

Each time around the circle will give another solution.

OpenStudy (mtalhahassan2):

don't you get 5pie/3

OpenStudy (mertsj):

What is 1/3 + 6/3 ?

OpenStudy (mtalhahassan2):